The expensive ones cost $7 each, whereas the worthless ones sold for only $2 each. You spent $111 and bought three more expensive ones than worthless ones. How many did you buy.
I know it is a system of equations: 7Ne + 2Nw = 111 is the first equation, don't know what the second equation is?

Answer by ReadingBoosters(3246)

Second equation is
Ne = Nw + 3
...
7(w+3) + 2w = 111
7w + 21 + 2w = 111
9w = 111 - 21
9w = 90
w = 10 worthless
...
e = 10 + 3 = 13 expensive
.....................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com